OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 32 + 8 ÷ 4 = 40 ÷ 4 = 10 B. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 C. 4 + 12 ÷ 2 = 4 + 6 = 10 D. (20 – 5) • 3 = 15 • 3 = 45

OpenStudy (anonymous):

32 + 8 ÷ 4 = 40 ÷ 4 = 10 B. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 C. 4 + 12 ÷ 2 = 4 + 6 = 10 D. (20 – 5) • 3 = 15 • 3 = 45

OpenStudy (anonymous):

The order of operations can be remembered with PEMDAS. From left-to-right, always operate in this order: Parentheses (including absolute-value signs, etc.), Exponents, Multiplication/Division, Addition/Subtraction. (A) We can see the violation in addition occurring first. FALSE (B) We can see the violation in the expression inside parentheses not being evaluated first. FALSE (C) Operation was done in the correct order. TRUE (D) Operation was done in the correct order. TRUE